# Generalized Virial Identities: Radial Constraints for Solitons, Instantons, and Bounces

### v2

@@ 1 Introduction
Topological solitons and tunneling configurations occupy a central place in modern field theory MantonSutcliffe2004; Shnir2005. Magnetic monopoles tHooft1974; Polyakov1974 arise as inevitable consequences of grand unified symmetry breaking and carry implications for early-universe cosmology. Cosmic strings and vortices NielsenOlesen1973; Vilenkin1985 seed density perturbations and produce distinctive gravitational wave signatures. The electroweak sphaleron KlinkhamerManton1984; MatchevVerner2025 mediates baryon number violation at high temperatures, while Skyrmions Skyrme1962; AdkinsNappiWitten1983; MantonSkyrmions2022 model baryons in the large-$N_{c}$ limit of QCD. Bounce solutions Coleman1977; ColemanCallan1977; ColemanDeLuccia1980 govern vacuum decay rates in scalar field theories and play an essential role in discussions of Higgs metastability. These configurations share a common mathematical structure: they extremize energy or action functionals subject to topological or boundary constraints, and their properties (mass, size, interaction potential) are encoded in the profile functions that describe them.
While a handful of these solutions admit closed-form expressions (the BPS monopole, the BPST instanton, the Fubini-Lipatov instanton), most must be obtained numerically. The question we address in this paper is not how to obtain such solutions, but rather how to characterize their radial structure through integral identities that encode the underlying physics. The most familiar such identity is Derrick’s theorem Derrick1964, which provides a single integral constraint relating total kinetic and potential contributions. Under a uniform dilation $x\to\lambda x$, these contributions scale differently, and stationarity at $\lambda=1$ yields a relation that any equilibrium configuration must satisfy. This relation has profound consequences: it forbids stable scalar solitons in three or more spatial dimensions when the potential is positive-definite, and it explains why the Skyrme term (quartic in derivatives) is necessary for Skyrmion stability. However, the Derrick relation is a global constraint that integrates over all radii and reveals nothing about the local balance between kinetic and potential contributions in different regions of the soliton.
@@
Our work connects to several recent developments in integral identities for solitons. Manton Manton2009 derived scaling identities beyond Derrick’s theorem by considering more general coordinate transformations, applying infinitesimal diffeomorphisms to the action functional. Our $\alpha$-family overlaps with Manton’s identities at specific parameter values but differs in emphasis: we parameterize the identities continuously and focus on their utility as radial probes. Adam et al. Adam2024 demonstrated that for any nonlinear field theory supporting static solutions with finite energy, an infinite number of integral identities can be derived from field transformations and their Noether currents. They showed that identities generated by coordinate transformations become trivial for BPS solitons, while field-space transformations yield nontrivial relations connected to observables such as the $D$-term and mechanical radius PolyakovSchweitzer2018; Ji2021. Our formalism provides explicit realizations of coordinate-based identities for O($n$) symmetric configurations and confirms the BPS triviality. Gudnason, Gao, and Yang GudnasonGaoYang2017 established a related 3-parameter family of integral identities for spherically symmetric solitons using boundary charge methods; our approach differs in deriving the identities from the radial functional structure. Herdeiro et al. HerdeiroOliveiraPomboRadu2021; HerdeiroOliveiraPomboRadu2022 extended virial identities to relativistic gravity, showing how they can be recast as energy-momentum balance conditions for self-gravitating solitons and black holes. The $\alpha$-family also generalizes the classical Pohozaev identities Pohozaev1965 from elliptic PDE theory: Pohozaev derived integral constraints by multiplying the equation $\Delta u=f(u)$ by $x\cdot\nabla u$ and integrating by parts, which for radial solutions corresponds to $\alpha=1$. Our generalization to arbitrary $\alpha$ yields a continuous family of constraints.
@@ 2.6 Derrick’s theorem
Setting $\alpha=1$ in Eq. (13) yields (2n)T=nU(2-n)T=nU, which is Derrick’s theorem Derrick1964. For $n\geq 3$ with $T,U>0$ (positive-definite kinetic energy and non-negative potential bounded below), this constraint cannot be satisfied: the kinetic and potential terms scale in the same direction under uniform dilation, and no equilibrium exists. This explains the absence of stable scalar solitons in three or more spatial dimensions.
@@ 3.1 Fubini-Lipatov instanton
The exact instanton solution, discovered independently by Fubini and Lipatov Rajaraman1982, takes the form
@@ 3.2 Vacuum decay
The decay of a metastable vacuum proceeds via quantum tunneling, described semiclassically by an O(4) symmetric bounce solution Coleman1977. The tunneling rate per unit volume takes the form $\Gamma/V\sim Ae^{-B}$, where $B$ is the Euclidean action of the bounce and $A$ is a prefactor arising from fluctuations around the classical solution. Because $B$ enters the exponential, even modest fractional errors in the bounce action translate into order-of-magnitude uncertainties in the decay rate.
@@
and $V$ is shifted so that $V(\phi_{\rm fv})=0$. The ratio $I_{T}^{\alpha}/I_{V}^{\alpha}=(\alpha+3)/(\alpha-3)$ takes specific values at each $\alpha$: 1-1 at $\alpha=0$, 2-2 at $\alpha=1$, and 5-5 at $\alpha=2$.
@@
Since $\dot{\phi}^{2}\geq 0$ and the bounce is nontrivial, we conclude that $B>0$. This positivity ensures exponential suppression of the decay rate $\Gamma\sim e^{-B}$, a result first established by Coleman Coleman1977 and Weinberg Weinberg1996.
@@
coincides with the mechanical energy in Coleman’s analogy  Coleman1977. The evolution equation (5) at $\alpha=-3$ becomes
@@
We computed the bounce numerically for the potential (33) with $\lambda=1$, $a=1$, and $\epsilon=0.1$. The resulting profile is shown in Figure 1. The bounce action for these parameters is $B\approx 24702$. Table 1 presents the numerical verification of the virial relations across a range of $\alpha$ values.
@@ 4.1 BPS monopole
The ’t Hooft-Polyakov monopole tHooft1974; Polyakov1974 offers a stringent test of the virial formalism in gauge theories. This solution arises in SU(2) Yang-Mills-Higgs theory with Lagrangian
@@
where $\Phi^{a}$ is a triplet Higgs field breaking SU(2) to U(1). In the BPS limit Bogomolny1976 $\lambda\to 0$, the monopole saturates a topological energy bound and admits an exact analytical solution.
@@
where $I_{\dot{K}}^{\alpha}=\int_{0}^{\infty}(\dot{K})^{2}\rho^{\alpha-1}\,d\rho$, etc. This is the hallmark of BPS configurations: the first-order equations guarantee automatic satisfaction of the virial identity for all valid $\alpha$.
@@
Using the BPS identities (55), both sides reduce to $\alpha I_{\dot{K}}^{\alpha}+(\alpha-2)I_{\dot{H}}^{\alpha}$, so the identity is satisfied tautologically for all valid $\alpha$. Since the monopole fields decay exponentially at large $\rho$ and behave as $K-1\sim\rho^{2}$, $H\sim\rho$ near the origin, the integrals converge for $\alpha>-2$.
@@
For the BPS monopole, the Bogomolny equations imply $\mathcal{C}_{\rho}=0$ pointwise, so the boundary term vanishes identically. However, the valid $\alpha$ range is still determined by integral convergence.
@@
At infinity ($\rho\to\infty$): The gauge field decays exponentially while the Higgs field approaches its vacuum algebraically:
$K\sim 2\rho\,e^{-\rho},\quad H\to 1-\frac{1}{\rho}.$ (59)
Since $\mathcal{C}_{\rho}=0$ exactly for the BPS solution, the boundary term vanishes identically. The remaining constraint comes from integral convergence at infinity.
@@ 4.2 BPST instanton
The BPST instanton BPST1975 in four-dimensional pure Yang-Mills theory offers another example where conformal invariance renders the standard Derrick test trivial, while the $\alpha\neq 1$ relations remain nontrivial. The Euclidean action is
@@
At $\alpha=1$, both sides vanish identically: the standard Derrick scaling argument provides no constraint because the action is scale-invariant. This is a trivial satisfaction of the $\alpha=1$ identity.
@@
which is a nontrivial constraint on the action density distribution. Both integrals scale as $\rho_{0}^{\alpha-1}$ under rescaling, so the identity is satisfied for any $\rho_{0}$ as required by conformal invariance, but the form of the profile is constrained.
@@
In this case the realtions hold for $3<\alpha<5$. Like the Fubini-Lipatov instanton, the finite upper bound reflects power-law decay in this conformally invariant theory.
@@ 4.3 Nielsen-Olesen vortices
The Nielsen-Olesen vortex NielsenOlesen1973 is the fundamental topological defect in the Abelian Higgs model with Lagrangian
@@ 4.4 Electroweak sphaleron
The electroweak sphaleron illustrates the virial formalism in a system where the Higgs mechanism explicitly breaks scale invariance, introducing fixed length scales absent in the pure Yang-Mills theory. In the Standard Model, the sphaleron is the static saddle-point solution at the top of the energy barrier separating topologically distinct vacua, playing a central role in baryon number violation in the early universe KlinkhamerManton1984; MatchevVerner2025.
@@
For the spherically symmetric ansatz with gauge profiles $f_{A}$, $f_{B}$ and Higgs profiles $H$, $K$ MatchevVerner2025, the energy functional becomes
@@ 5 Skyrmions
The Skyrme model provides an effective field theory of pions whose topological solitons, Skyrmions, model baryons in the large-$N_{c}$ limit of QCD Skyrme1962; AdkinsNappiWitten1983. The Lagrangian is
@@
Unlike the gauge-Higgs systems of the previous section, the Skyrme model requires a term quartic in derivatives for soliton stability. Derrick’s theorem forbids static solitons when all terms scale identically under dilation; the Skyrme term, with its different scaling dimension, provides the resistance to collapse that allows equilibrium. Manton Manton2024 recently demonstrated that the hedgehog profile varies only weakly across different chiral effective field theories. The virial identities developed here provide a systematic understanding of this robustness: at special $\alpha$ values, different terms decouple, revealing which contributions dominate in different radial regions.
@@ 5.2 Standard Skyrme model
The virial identity (106) simplifies at particular $\alpha$ values, isolating specific balances:
@@
Manton notes that $K_{2,0,0}$, $K_{2,2,0}$, and $K_{2,2,2}$ are the only symmetric polynomials quadratic in time-derivatives Manton2024. The sextic contribution $K_{2,2,2}=3\dot{f}^{2}\sin^{4}f/\rho^{4}$ has coefficient $\alpha+2$ in the virial identity, yielding
@@ 5.3 Asymptotic universality
This confirms Manton’s observation Manton2024 that the tail is determined entirely by the Dirichlet term, with the angular-strain terms becoming negligible at large $\rho$. The virial formalism provides a systematic derivation of this asymptotic universality.

### v3

@@ 1 Introduction
Topological solitons and tunneling configurations occupy a central place in modern field theory Manton and Sutcliffe (2004); Shnir (2005). Magnetic monopoles ’t Hooft (1974); Polyakov (1974) arise as inevitable consequences of grand unified symmetry breaking and carry implications for early-universe cosmology. Cosmic strings and vortices Nielsen and Olesen (1973); Vilenkin (1985) seed density perturbations and produce distinctive gravitational wave signatures. The electroweak sphaleron Klinkhamer and Manton (1984a); Matchev and Verner (2025) mediates baryon number violation at high temperatures, while Skyrmions Skyrme (1962); Adkins et al. (1983); Manton (2022) model baryons in the large-$N_{c}$ limit of QCD. Bounce solutions Coleman (1977); Callan and Coleman (1977); Coleman and De Luccia (1980) govern vacuum decay rates in scalar field theories and play an essential role in discussions of Higgs metastability. These configurations share a common mathematical structure: they extremize energy or action functionals subject to topological or boundary constraints, and their properties (mass, size, interaction potential) are encoded in the profile functions that describe them.
While a handful of these solutions admit closed-form expressions (the BPS monopole, the BPST instanton, the Fubini-Lipatov instanton), most must be obtained numerically. The question we address in this paper is not how to obtain such solutions, but rather how to characterize their radial structure through integral identities that encode the underlying physics. The most familiar such identity is Derrick’s theorem Derrick (1964), which provides a single integral constraint relating total kinetic and potential contributions. Under a uniform dilation $x\to\lambda x$, these contributions scale differently, and stationarity at $\lambda=1$ yields a relation that any equilibrium configuration must satisfy. This relation has profound consequences: it forbids stable scalar solitons in three or more spatial dimensions when the potential is positive-definite, and it explains why the Skyrme term (quartic in derivatives) is necessary for Skyrmion stability. However, the Derrick relation is a global constraint that integrates over all radii and reveals nothing about the local balance between kinetic and potential contributions in different regions of the soliton.
@@
Our work connects to several recent developments in integral identities for solitons. Manton Manton (2009) derived scaling identities beyond Derrick’s theorem by considering more general coordinate transformations, applying infinitesimal diffeomorphisms to the action functional. Our $\alpha$-family overlaps with Manton’s identities at specific parameter values but differs in emphasis: we parameterize the identities continuously and focus on their utility as radial probes. Adam et al. Adam et al. (2024) demonstrated that for any nonlinear field theory supporting static solutions with finite energy, an infinite number of integral identities can be derived from field transformations and their Noether currents. They showed that identities generated by coordinate transformations become trivial for BPS solitons, while field-space transformations yield nontrivial relations connected to observables such as the $D$-term and mechanical radius Polyakov and Schweitzer (2018); Ji (2021). Our formalism provides explicit realizations of coordinate-based identities for O($n$) symmetric configurations and confirms the BPS triviality. Gudnason, Gao, and Yang Gudnason et al. (2017) established a related 3-parameter family of integral identities for spherically symmetric solitons using boundary charge methods; our approach differs in deriving the identities from the radial functional structure. Herdeiro et al. Herdeiro et al. (2021, 2022) extended virial identities to relativistic gravity, showing how they can be recast as energy-momentum balance conditions for self-gravitating solitons and black holes. The $\alpha$-family also generalizes the classical Pohozaev identities Pohozaev (1965) from elliptic PDE theory: Pohozaev derived integral constraints by multiplying the equation $\Delta u=f(u)$ by $x\cdot\nabla u$ and integrating by parts, which for radial solutions corresponds to $\alpha=1$. Our generalization to arbitrary $\alpha$ yields a continuous family of constraints.
@@ 2.6 Derrick’s theorem
Setting $\alpha=1$ in Eq. (13) yields (2n)T=nU(2-n)T=nU, which is Derrick’s theorem Derrick (1964). For $n\geq 3$ with $T,U>0$ (positive-definite kinetic energy and non-negative potential bounded below), this constraint cannot be satisfied: the kinetic and potential terms scale in the same direction under uniform dilation, and no equilibrium exists. This explains the absence of stable scalar solitons in three or more spatial dimensions.
@@ 3.1 Fubini-Lipatov instanton
The exact instanton solution, discovered independently by Fubini and Lipatov Rajaraman (1982), takes the form
@@ 3.2 Vacuum decay
The decay of a metastable vacuum proceeds via quantum tunneling, described semiclassically by an O(4) symmetric bounce solution Coleman (1977). The tunneling rate per unit volume takes the form $\Gamma/V\sim Ae^{-B}$, where $B$ is the Euclidean action of the bounce and $A$ is a prefactor arising from fluctuations around the classical solution. Because $B$ enters the exponential, even modest fractional errors in the bounce action translate into order-of-magnitude uncertainties in the decay rate.
@@
and $V$ is shifted so that $V(\phi_{\rm fv})=0$. The normalization $V(\phi_{\mathrm{fv}})=0$ is required by the vanishing of the boundary term. On a finite interval $[0,R]$, the identity acquires a boundary contribution $R^{\alpha+n-1}(\frac{1}{2}\dot{\phi}(R)^{2}-V(\phi(R)))$; for solutions with infinite support this must vanish as $R\to\infty$, while for compactons with $\dot{\phi}(R)=0$ it vanishes at finite $R$. Both cases require $V(\phi_{\mathrm{fv}})=0$. The ratio $I_{T}^{\alpha}/I_{V}^{\alpha}=(\alpha+3)/(\alpha-3)$ takes specific values at each $\alpha$: 1-1 at $\alpha=0$, 2-2 at $\alpha=1$, and 5-5 at $\alpha=2$.
@@
Since $\dot{\phi}^{2}\geq 0$ and the bounce is nontrivial, we conclude that $B>0$. This positivity ensures exponential suppression of the decay rate $\Gamma\sim e^{-B}$, a result first established by Coleman Coleman (1977) and Weinberg Weinberg (1996).
@@
coincides with the mechanical energy in Coleman’s analogy  Coleman (1977). The evolution equation (5) at $\alpha=-3$ becomes
@@
We computed the bounce numerically for the potential (3.18) with $\lambda=1$, $a=1$, and $\epsilon=0.1$ using the AnyBubble package Masoumi et al. (2017). The resulting profile is shown in Figure 1. The bounce action for these parameters is $B\approx 24725.2$. Table 1 presents the numerical verification of the virial relations across a range of $\alpha$ values.
@@ 4.1 BPS monopole
The ’t Hooft-Polyakov monopole ’t Hooft (1974); Polyakov (1974) offers a stringent test of the virial formalism in gauge theories. This solution arises in the Georgi–Glashow model Georgi and Glashow (1972), i.e.,SU(2) Yang-Mills-Higgs theory with Lagrangian
@@
where $\Phi^{a}$ is a triplet Higgs field breaking SU(2) to U(1). In the BPS limit Bogomolny (1976) $\lambda\to 0$, the monopole saturates a topological energy bound and admits an exact analytical solution.
@@
where $I_{\dot{K}}^{\alpha}=\int_{0}^{\infty}(\dot{K})^{2}\rho^{\alpha-1}\,d\rho$, etc. This is the hallmark of BPS configurations: the first-order BPS equations are equivalent to the condition of vanishing pressure, from which the trivial satisfaction of all integral relations follows immediately.
@@
Using the BPS identities (55), both sides reduce to $\alpha I_{\dot{K}}^{\alpha}+(\alpha-2)I_{\dot{H}}^{\alpha}$, so the identity is satisfied tautologically for all valid $\alpha$.
@@
For the BPS monopole, the Bogomolny equations imply $\mathcal{C}_{\rho}=0$ pointwise, so the boundary term vanishes identically.
@@
At infinity ($\rho\to\infty$): The unbroken $U(1)$ symmetry produces a massless photon whose Coulomb field gives power-law tails in both the gauge and Higgs sectors. From the exact solution (53), the gauge profile decays exponentially, $K\sim 2\rho\,e^{-\rho}$, while the Higgs profile approaches its VEV algebraically, $H-1\sim-1/\rho$. Of the four integrals in (56), $I_{\dot{H}}^{\alpha}$ and $I_{K^{2}}^{\alpha}$ have integrands falling as $\rho^{\alpha-3}$ and diverge individually for $\alpha\geq 2$, while $I_{\dot{K}}^{\alpha}$ and $I_{KH}^{\alpha}$ decay exponentially. The Bogomolny equations enforce exact pairwise cancellations: $I_{\dot{K}}^{\alpha}=I_{KH}^{\alpha}$ and $I_{\dot{H}}^{\alpha}=I_{K^{2}}^{\alpha}$, so the virial identity reduces to $0=0$ for all $\alpha$. No upper bound on $\alpha$ arises from infinity.
@@ 4.2 BPST instanton
The BPST instanton Belavin et al. (1975) in four-dimensional pure Yang-Mills theory provides a second example of BPS triviality, where the self-duality equation $F=\ast F$ guarantees that the virial identities are satisfied for all valid $\alpha$. Additionally, conformal invariance renders the $\alpha=1$ identity degenerate. The Euclidean action is
@@
The self-duality equation $F=\ast F$ reduces in the radial problem to
$\rho\dot{w}=2w(1-w).$ (64)
Squaring both sides gives $\rho^{2}\dot{w}^{2}=4w^{2}(1-w)^{2}$, i.e., the kinetic and potential densities in (61) are related pointwise:
$\frac{1}{2}\rho\dot{w}^{2}=\frac{2}{\rho}w^{2}(1-w)^{2}.$ (65)
This implies $\mathcal{C}_{\rho}=0$ exactly, just as for the BPS monopole. The virial identities are therefore trivially satisfied for all valid $\alpha$.
@@
At $\alpha=1$, both sides vanish identically: the standard Derrick scaling argument provides no constraint because the action is scale-invariant. This degeneracy is specific to conformally invariant theories and is independent of the BPS condition.
@@
which is the self-duality equation squared and integrated with weight $\rho^{\alpha-2}$. Both integrals scale as $\rho_{0}^{\alpha-1}$ under rescaling, so the identity is satisfied for any $\rho_{0}$ as required by conformal invariance.
@@
In this case the relations hold for $3<\alpha<5$. Like the Fubini-Lipatov instanton, the finite upper bound reflects power-law decay in this conformally invariant theory.
@@ 4.3 Nielsen-Olesen vortices
The Nielsen-Olesen vortex Nielsen and Olesen (1973) is the fundamental topological defect in the Abelian Higgs model with Lagrangian
@@ 4.4 Electroweak sphaleron
The electroweak sphaleron illustrates the virial formalism in a system where the Higgs mechanism explicitly breaks scale invariance, introducing fixed length scales absent in the pure Yang-Mills theory. In the Standard Model, the sphaleron is the static, unstable saddle-point solution of the classical field equations Klinkhamer and Manton (1984b); Manton (2019), situated at the energy barrier separating topologically distinct vacua playing a central role in baryon number violation in the early universe Klinkhamer and Manton (1984a); Matchev and Verner (2025).
@@
For the spherically symmetric ansatz with gauge profiles $f_{A}$, $f_{B}$ and Higgs profiles $H$, $K$ Matchev and Verner (2025), the energy functional becomes
@@ 5 Skyrmions
The Skyrme model provides an effective field theory of pions whose topological solitons, Skyrmions, model baryons in the large-$N_{c}$ limit of QCD Skyrme (1962); Adkins et al. (1983). The Lagrangian is
@@
Unlike the gauge-Higgs systems of the previous section, the Skyrme model requires a term quartic in derivatives for soliton stability. Derrick’s theorem forbids static solitons when all terms scale identically under dilation; the Skyrme term, with its different scaling dimension, provides the resistance to collapse that allows equilibrium. Manton Manton (2024) recently demonstrated that the hedgehog profile varies only weakly across different chiral effective field theories. The virial identities developed here provide a systematic understanding of this robustness: at special $\alpha$ values, different terms decouple, revealing which contributions dominate in different radial regions.
@@ 5.2 Standard Skyrme model
For the standard Skyrme model, the identity (107) was first obtained by Gudnason, Gao, and Yang Gudnason et al. (2017) (their Eq. (4.30) with $m=0$ and $2\kappa=\alpha+2$ by using . We rederive it here within our formalism and discuss its special $\alpha$ values. The virial identity (107) simplifies at particular $\alpha$ values, isolating specific balances:
@@
Manton notes that $K_{2,0,0}$, $K_{2,2,0}$, and $K_{2,2,2}$ are the only symmetric polynomials quadratic in time-derivatives Manton (2024). The sextic contribution $K_{2,2,2}=3\dot{f}^{2}\sin^{4}f/\rho^{4}$ has coefficient $\alpha+2$ in the virial identity, yielding
@@ 5.3 Asymptotic universality
This confirms Manton’s observation Manton (2024) that the tail is determined entirely by the Dirichlet term, with the angular-strain terms becoming negligible at large $\rho$. The virial formalism provides a systematic derivation of this asymptotic universality.